# Evolving Agents: Interactive Simulation of Dynamic and Diverse Human Personalities

### v1

@@ Abstract.
Human-like Agents with diverse and dynamic personality could serve as an important design probe in the process of user-centered design, thereby enabling designers to enhance the user experience of interactive application.
In this article, we introduce Evolving Agents, a novel agent architecture that consists of two systems: Personality and Behavior. The Personality system includes three modules: Cognition, Emotion and Character Growth. The Behavior system comprises two modules: Planning and Action. We also build a simulation platform that enables agents to interact with the environment and other agents.
Evolving Agents can simulate the human personality evolution process. Compared to its initial state, agents’ personality and behavior patterns undergo believable development after several days of simulation. Agents reflect on their behavior to reason and develop new personality traits. These traits, in turn, generate new behavior patterns, forming a feedback loop-like personality evolution.
In our experiment, we utilized simulation platform with 10 agents for evaluation. During the evaluation, these agents experienced believable and inspirational personality evolution. Through ablation and control experiments, we demonstrated the outstanding effectiveness of agent personality evolution and all modules of our agent architecture contribute to creating believable human-like agents with diverse and dynamic personalities. We also demonstrated through workshops how Evolving Agents could inspire designers.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: xxx; ; ccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Natural language processing
@@
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
@@ 1. INTRODUCTION
Embodied agents are capable of simulating the humans behaviors (kampmann_feedback_2012, ). By observing and interacting with these human-like agents, we can gain a better understanding of ourselves as humans. Within a design context, analyzing how these agents operate in specific scenarios allows designers to uncover insights into potential user behaviors, needs, and experiences, making the embodied agents an inspirational design probe in the user-centered design process.
@@
Numerous studies have already proposed embodied agents’ capable of simulating human-like behaviors within virtual environments. Agents can actively interact with objects and other agents in the simulation environment(noauthor_generative_nodate, )(wang_humanoid_2023, )(lin_agentsims_2023, ) or continuously learn new behaviors guided by predefined tasks in the environment(wang_voyager_2023, )(fan_minedojo_2022, ).
@@
One of the foremost objectives in the development of embodied agents is to enhance their behavior’s similarity to human actions within the simulation environment(zhu_ghost_2023, )(chen_llm-empowered_2023, )(du_improving_2023, )(zhang_building_2024, ). Real humans exhibit a wide array of personality traits and corresponding behavior patterns(zuckerman_psychobiology_1991, ) that dynamically evolve through interacting with the external world(ghasem-aghaee_cognitive_2007, ). However, current embodied agents show a constrained range of personality expressions and, notably, their behaviors and personalities do not undergo evolution over time via continuous interaction with the simulation environments. In previous sandbox simulation studies, there are seldom instances of the personality evolution process occurring without task-oriented guidance.Agents that do not undergo continuous evolution seem to experience each day as one long moment, repeating itself over and over.
@@
Furthering this point, we find that current agents rarely exhibit diversified and dynamic personalities within the interactive simulation process. This results in a lack of variation in behavioral patterns both among different agents and across different simulation periods of the same agent. Therefore, we aim to design a novel agent architecture that enables agents to possess their own diversified and dynamic personalities, that could make agents experience their days as thousands of single and different moments, which is called personality evolution in previous work.(oren_dynamic_1991, )
@@
To realize agents personality evolution, enabling agents to exhibit differentiated behaviors corresponding to their various personalities and undergo continuous personality evolution based on external information, during the interactive simulation system, we drew inspiration from psychological theories and proposed a novel architecture - Evolving Agents. Through simulation and experimental analysis in the sandbox environments, Evolving Agents are capable of simulating believable personality traits and personalized behavioral patterns. Moreover, their personalities and behaviors undergo continuous evolution throughout the simulation process. Evolving Agents exhibit differentiated behavioral patterns based on their diverse personality, while their behavioral experiences also influence the changes in agent personalities at the same time, forming a feedback loop-like process of behavior-personality development. For example, Benjamin, committed to changing his Introverted personality, so he makes efforts to engage in more interpersonal communication. His enjoyable social experiences further enhance Benjamin’s extroversion, ultimately leading to a more open-minded attitude towards interpersonal interactions. Similarly, Isabella actively explores her personal values during the process, and the encouragement she receives from others reinforces her determination to live for herself, consequently enhancing Isabella’s self-awareness.
@@
The architectural design of Evolving Agents incorporates insights from both psychology and computer science. By adopting the stimulus-response concept from behavioral psychology, we have developed two main systems: the Behavior System, which is responsible for planning specific behaviors and execution. The Personality system, constructed with Emotion, Cognition, and Character Growth modules, cognitively processes external information to enable continuous evolution. The interaction between these two systems allows agents to exhibit human-like personalized behavior patterns and undergo dynamic personality evolution. Additionally, we have developed a new sandbox environment where agents can interact with environmental scenarios and other agents. This platform allows users to edit environments and interact with agents, providing intuitive and engaging simulations.
@@
To validate the perception and believability of Evolving Agents simulating human diverse personality traits and dynamic personality evolution, we proposed several research hypotheses:
@@
We conducted both objective data analysis and subjective human evaluation, with both sets of results preliminarily confirming our research hypotheses. In the objective data analysis, we utilized the Big Five personality traits scale and behavioral similarity analysis to verify whether Evolving Agents have undergone differentiated and dynamic personality evolution, and whether the behavior pattern develops continuously. In human evaluation, we engaged 30 human assessors to evaluate the believability of personality differences and the personality evolution in Evolving Agents. They were also asked to compare several ablations that restrict the agents’ access to Character Growth module and complete Character Structure. These demonstrates that Evolving Agents can perceptibly and believably exhibit differentiated behaviors in alignment with their personalities, as well as undergo continuous and believable personality evolution, and each of these components is critical to strong performance.Furthermore, we organized designer workshops around this system to discuss the inspiring effect of Evolving Agents as design probes for designers, yielding valuable insights into integrating such systems into user-oriented design processes.
@@
  • A novel agent architecture, having behavior and personality two main systems, makes it possible to execute planning, action, emotion, cognition and character growth in the Virtual Simulation Environment, enabling agents to exhibit differentiated behaviors corresponding to their various personalities and undergo continuous personality evolution based on external information during the interactive simulation.
  • An evaluation procedure was established that includes two methods: objective data verification and human evaluation, for assessing the effectiveness of agent personality evolution and the contributions of each module within the architecture during this process. Additionally, a supportive workshop was conducted to discuss the inspirational effects of Evolving Agents on designers.
@@ 2.1. Interaction Between Humans And Agents
As generative models advance rapidly in general capabilities and intent recognition(zuckerman_psychobiology_1991, )(hamilton_blind_2023, )(swan_math_2023, ), agents powered by these models have found widespread application across various fields and task types(xi_rise_2023, )(gur_real-world_2024, )(wu_plan_2023, ), demonstrating a broad spectrum of application prospects. With the enrichment of application scenarios for agents, novel patterns of interaction between humans and agents are continually emerging.
Human-agent interaction, as the name suggests, involves agents collaborating with humans to accomplish tasks(xi_rise_2023, ). Such as introducing agents to assist in peer psychological counseling(hsu_helping_2023, ), using agents in telemarketing scenarios(zamfirescu-pereira_why_2023, ), or co-creating literature or artworks with agents(zamfirescu-pereira_why_2023, )(bhat_interacting_2023, )(noauthor_3dall-e_nodate, ).
In recent years, the human-like behavior exhibited by embodied agents in simulation environments has attracted widespread social attention(kalvakurthi_hey_2023, )(hsu_helping_2023, )(hasan_sapien_2023, )(noauthor_human-level_nodate, ). Among these, the Generative Agent discussed human social and communication behaviors within societies. By observing simulated behaviors of agents in a sandbox environment(noauthor_generative_nodate, ), we can know ourselves as humans better.
Some research focuses on multi-agent simulation in more goal-oriented settings, such as collaborative efforts in software production within software studios(li_camel_2023, ). While some others focused on creating agents capable of autonomous exploration and continuous learning within game worlds(wang_voyager_2023, ). Additionally, there are studies aimed at developing human-like NPCs, enabling text-based strategy game interactions with users, provide imaginative possibilities for future interaction possibilities between game players and NPCs or companions.
When focusing on the interaction between designers and agents, designers can observe the simulated behaviors and feedback of agents representing target users, gaining valuable design analysis materials to assist in the user-oriented design process(zamfirescu-pereira_why_2023, )(zamfirescu-pereira_herding_2023, ). Current research has explored this direction; for example, one study aimed to simulate doctor-patient communication scenarios in psychological counseling to provide references and suggestions for real treatment(chen_llm-empowered_2023, ). Another study simulated user behaviors in scenarios like online movie ratings, offering insights for further exploration in these areas(ren_bases_2024, ).
@@ 2.2. Human-like Agents In Virtual Simulation Environment
Simulation of human behavior in real-world like environment has been an interesting and important area for a long time(park_generative_2023, ).Ushered in by the landmark paper on Generative Agents(park_generative_2023, ), the promise of modelling conceivable human-like behavior in Virtual Simulation Environment has sparked the imagination of many(wang_survey_2024, ).
@@
The human-like dimension of embodied agents can be subdivided into intrinsic and extrinsic aspects(noauthor_becoming_nodate, )(xi_rise_2023, ), as informed by sociological opinions. The extrinsic aspect encompasses behaviors at both the individual and group levels, whereas the intrinsic aspect comprises emotions, characteristics, and cognition. These delineated dimensions afford researchers numerous opportunities to enhance the human-likeness of agents(yao_react_2023, )(chen_llm-empowered_2023, )(driess_palm-e_2023, )(lin_agentsims_2023, )(park_generative_2023, ).
@@
Current research on human-like agents in simulation environments predominantly focuses on their extrinsic aspects. At the level of individual behavior, Danny Driess(serapio-garcia_personality_2023, )(du_improving_2023, )(wu_autogen_2023, ) proposed a multi-modal task-executing agent designed for real-world scenarios. Projects such as MINEDOJO and Voyager, leveraging the expansive game world of Minecraft, explore the individual learning behaviors of agents within such virtual settings(lin_agentsims_2023, )(noauthor_230516291_nodate, )(fan_minedojo_2022, ).
@@
Regarding group-level behaviors, Generative Agents have been developed to enable agents to perform social actions. Additionally, AgentSim offers a comprehensive architecture for simulating standardized agent behaviors, incorporating a module for tool use among agents.
@@
Nowadays, Researchers are gradually realizing the practical computational significance of an Agent’s personality on human-like behaviors(ghasem-aghaee_cognitive_2007, )(liu_training_2023, )(kalvakurthi_hey_2023, )(varela_embodied_2017, ). In the study of Humanoid Agents(wang_humanoid_2023, ), researchers explored how emotions and needs influence an Agent’s planning, largely addressing the overly rational and stable machine-like state exhibited by former Agents. However, research on embodied agents in virtual environments seldom discuss other intrinsic dimensions. Expanding the scope to other Agent application scenarios, such as conversational interaction Chatbots(tu_characterchat_2023, ), we can still find some works focusing on the intrinsic aspects of Agents, such as employing psychological theories, especially cognitive psychology, to test and enhance the cognitive abilities of personality agents(habibi_empathetic_2023, ), and some works focusing on modeling emotions, especially emotional recognition and empathy(habibi_empathetic_2023, ). In terms of character, some works simulate the static personality and knowledge of historical figures based on external data(yao_react_2023, ), and others create chatbots with different static personalities based on personality psychology theories(tu_characterchat_2023, ).
@@
It should be noted that the intrinsic personality of real humans is not immutable. In the field of psychology, researchers believe that personality is continuously influenced and changed by external factors(kampmann_feedback_2012, ). In existing discussions on research involving embodied agents, we rarely find studies that detail the process of such human-like intrinsic personality evolution or studies on how intrinsic personality evolution affects an agent’s extrinsic behavior.

### v3

@@ Abstract.
Human-like Agents with diverse and dynamic personalities could serve as an important design probe in the user-centred design process, thereby enabling designers to enhance the user experience of interactive applications.
In this article, we introduce Evolving Agents, a novel agent architecture that consists of two systems: Personality and Behavior. The Personality system includes Cognition, Emotion, and Character Growth modules. The Behavior system comprises two modules: Planning and Action. We also build a simulation platform that enables agents to interact with the environment and other agents.
Evolving Agents can simulate the human personality evolution process. Compared to its initial state, agents’ personalities and behavior patterns undergo believable development after several days of simulation. Agents reflect on their behavior to reason and develop new personality traits. These traits, in turn, generate new behavior patterns, forming a feedback loop-like personality evolution.
Our experiment utilized a simulation platform with ten agents for evaluation. During the evaluation, these agents experienced believable and inspirational personality evolution. We demonstrated the effectiveness of agent personality evolution through ablation and control experiments. All of our agent architecture modules contribute to creating believable human-like agents with diverse and dynamic personalities. We also demonstrated through workshops how Evolving Agents could inspire designers.
@@
copyright: acmlicensedjournalyear: 2024doi: 2404.02718conference: xxxx; xxxxx; xxx, xxxxccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Natural language processing
@@
Enjoying the baseball game from the third-base seats. Ichiro Suzuki is preparing to bat.
@@ 1. INTRODUCTION
Embodied agents are capable of simulating the human’s behaviors. By observing and interacting with these human-like agents, we can better understand ourselves as humans. Within a design context, analyzing how these agents operate in specific scenarios allows designers to uncover insights into potential user behaviors, needs, and experiences(hong_conceptualization_2017, ; mcginn_datadriven_2008, ), making the embodied agents an inspirational design probe in the user-centred design process.
@@
Numerous studies have proposed that embodied agents can simulate human-like behaviors within virtual environments. Agents can actively interact with objects and other agents in the simulation environment(noauthor_generative_nodate, ; wang_humanoid_2023, ; lin_agentsims_2023, ) or continuously learn new behaviors guided by predefined tasks in the environment(wang_voyager_2023, ; fan_minedojo_2022, ).
@@
One of the foremost objectives in the development of embodied agents is to enhance their behavior’s similarity to human actions within the simulation environment(zhu_ghost_2023, ; chen_llm-empowered_2023, ; du_improving_2023, ; zhang_building_2024, ). Real humans exhibit a wide array of personality traits and corresponding behavior patterns(zuckerman_psychobiology_1991, ) that dynamically evolve through interacting with the external world(ghasem-aghaee_cognitive_2007, ). However, current embodied agents show a constrained range of personality expressions. However, their behaviors and personalities do not undergo evolution over time via continuous interaction with the simulation environments. In previous sandbox simulation studies, there are seldom instances of the personality evolution process occurring without task-oriented guidance. Agents that do not undergo continuous evolution seem to experience each day as one long moment, repeating itself repeatedly.
@@
Furthering this point, current agents rarely exhibit diversified and dynamic personalities within the interactive simulation process. This results in a lack of variation in behavioral patterns both among different agents and across different simulation periods of the same Agent. Therefore, we aim to design a novel agent architecture that enables agents to possess their own diversified and dynamic personalities, which could make agents experience their days as thousands of single and different moments, which is called personality evolution in previous work(oren_dynamic_1991, ).
@@
To realize agents’ personality evolution, enabling agents to exhibit differentiated behaviors corresponding to their various personalities and undergo continuous personality evolution based on external information, we drew inspiration from psychological theories(lachman_cognitive_2016, ; kihlstrom_cognitive_2018, ; allport_pattern_1961, ; baars_cognitive_1993, ). We proposed a novel architecture - Evolving Agents. Evolving Agents can simulate believable personality traits and personalized behavioral patterns through simulation and experimental analysis in sandbox environments. Moreover, their personalities and behaviors undergo continuous evolution throughout the simulation process. Evolving Agents exhibit differentiated behavioral patterns based on their diverse personalities. At the same time, their behavioral experiences also influence the changes in agent personalities simultaneously, forming a feedback loop-like process of behavior-personality development(brun_engineering_2009, ). For example, Benjamin is committed to changing his Introverted personality and tries to engage in more interpersonal communication. His enjoyable social experiences further enhance Benjamin’s extroversion, leading to a more open-minded attitude toward interpersonal interactions. Similarly, Isabella actively explores her values during the process, and the encouragement she receives from others reinforces her determination to live for herself, consequently enhancing Isabella’s self-awareness.
@@
Evolving Agents’ architectural design incorporates insights from psychology and computer science. By adopting the stimulus-response concept from behavioral psychology(clark_learning_2018, ; woodward_discovery_1982, ), we have developed two main systems: the Behavior System, responsible for planning specific behaviors and execution. The Personality system, constructed with Emotion, Cognition, and Character Growth modules, cognitively processes external information to enable continuous evolution. The interaction between these two systems allows agents to exhibit human-like personalized behavior patterns and undergo dynamic personality evolution. Additionally, we have developed a new sandbox environment where agents can interact with environmental scenarios and other agents. This platform allows users to edit environments and interact with agents, providing intuitive and engaging simulations.
@@
To validate the perception and believability of Evolving Agents simulating diverse human personality traits and dynamic personality evolution, we proposed several research hypotheses:
@@
We conducted objective data analysis and subjective human evaluation, with both results preliminarily confirming our research hypotheses. In the objective data analysis, we utilized the Big Five personality traits scale and behavioral similarity analysis to verify whether Evolving Agents have undergone differentiated and dynamic personality evolution and whether the behavior pattern develops continuously. In human evaluation, we engaged 30 human assessors to evaluate the believability of personality differences and the personality evolution in Evolving Agents. They were also asked to compare several ablations restricting the agents’ access to the Character Growth module and complete Character Structure. These demonstrate that Evolving Agents can perceptibly and believably exhibit differentiated behaviors in alignment with their personalities and undergo continuous and believable personality evolution. Each of these components is critical to strong performance. Furthermore, we organized designer workshops around this system to discuss the inspiring effect of Evolving Agents as design probes for designers, yielding valuable insights into integrating such systems into user-oriented design processes.
@@
  • A novel agent architecture, having behavior and personality as two main systems, makes it possible to execute planning, Action, emotion, cognition, and character growth in the Virtual Simulation Environment, enabling agents to exhibit differentiated behaviors corresponding to their various personalities and undergo continuous personality evolution based on external information during the interactive simulation.
  • An evaluation procedure was established that includes two methods, objective data verification, and human evaluation, for assessing the effectiveness of agent personality evolution and the contributions of each module within the architecture during this process. Additionally, we conducted a supportive workshop to discuss the inspirational effects of Evolving Agents on designers.
@@ 2.1. Interaction Between Humans And Agents
As generative models advance rapidly in general capabilities and intent recognition(zuckerman_psychobiology_1991, ; hamilton_blind_2023, ; swan_math_2023, ; wei_emergent_2022, ; ouyang_training_2022, ), agents powered by these models have found widespread application across various fields and task types(xi_rise_2023, ; wei_emergent_2022, ; sumers_cognitive_2024, ; liu_training_2023, ), demonstrating a broad spectrum of application prospects. With the enrichment of application scenarios for agents, novel patterns of interaction between humans and agents are continually emerging(kim_one_2023, ; li_map_2024, ; wang_human-human_2020, ).
Human-agent interaction, as the name suggests, involves agents collaborating with humans to accomplish tasks(xi_rise_2023, ), such as introducing agents to assist in peer psychological counseling(hsu_helping_2023, ), using agents in telemarketing scenarios(noauthor_multi-turn_nodate, ), or co-creating literature or artworks with agents(bhat_interacting_2023, ; noauthor_3dall-e_nodate, ; chiou_designing_2023, ).
In recent years, the human-like behavior exhibited by embodied agents in simulation environments has attracted widespread social attention(kalvakurthi_hey_2023, ; hsu_helping_2023, ; hasan_sapien_2023, ; noauthor_human-level_nodate, ; schick_peer_2023, ). Among these, the Generative Agent discussed human social and communication behaviors within societies. By observing simulated behaviors of agents in a sandbox environment(noauthor_generative_nodate, ), we can know ourselves as humans better.
Some research focuses on multi-agent simulation in more goal-oriented settings, such as collaborative efforts in software production within software studios(qian_communicative_2023-1, ). At the same time, some others focused on creating agents capable of autonomous exploration and continuous learning within game worlds(wang_voyager_2023, ). Additionally, studies aim to develop human-like NPCs, enabling text-based strategy game interactions with users and providing imaginative possibilities for future interaction between game players and NPCs or companions.
When focusing on the interaction between designers and agents, designers can observe the simulated behaviors and feedback of agents representing target users, gaining valuable design analysis materials to assist in the user-oriented design process(zamfirescu-pereira_why_2023, ; zamfirescu-pereira_herding_2023, ). Current research has explored this direction; for example, one study aimed to simulate doctor-patient communication scenarios in psychological counselling to provide references and suggestions for real treatment(chen_llm-empowered_2023, ). Another study simulated user behaviors in scenarios like online movie ratings, offering insights for further exploration in these areas(hong_conceptualization_2017, ).
@@ 2.2. Human-like Agents In Virtual Simulation Environment
Simulation of human behavior in the real-world environment has been an interesting and important area for a long time(noauthor_computer_nodate, ; cote_textworld_2019, ; hausknecht_interactive_2020, ; wang_describe_2023, ; nottingham_embodied_2023, ). Ushered in by the landmark paper on Generative Agents(park_generative_2023, ), the promise of modelling conceivable human-like behavior in Virtual Simulation Environment has sparked the imagination of many(noauthor_230812503_nodate, ; li_hitchhikers_2023, ; wang_scienceworld_2022, ).
@@
As informed by sociological opinions, the human-like dimension of embodied agents can be subdivided into intrinsic and extrinsic aspects(noauthor_becoming_nodate, ; xi_rise_2023, ). The extrinsic aspect encompasses behaviors at both the individual and group levels(liu_llmp_2023, ; qian_communicative_2023-1, ; chan_chateval_2023, ; chen_agentverse_2023, ). In contrast, the intrinsic aspect comprises emotions, characteristics, and cognition(tu_characterchat_2023, ; jiang_evaluating_2023, ; li_large_2023, ; serapio-garcia_personality_2023, ; dhingra_mind_2023, ). These delineated dimensions afford researchers numerous opportunities to enhance the human-likeness of agents(yao_react_2023, ; hagendorff_machine_2023, ; driess_palm-e_2023, ; noauthor_frontiers_nodate, ; liu_training_2023, ; fan_minedojo_2022, ).
@@
Current research on human-like agents in simulation environments predominantly focuses on their extrinsic aspects. At the level of individual behavior, Danny Driess(driess_palm-e_2023, ) proposed a multi-modal task-executing agent designed for real-world scenarios. Projects such as MINEDOJO(fan_minedojo_2022, ) and Voyager(wang_voyager_2023, ), leveraging the expansive game world of Minecraft, explore the individual learning behaviors of agents within such virtual settings.
@@
Regarding group-level behaviors, Generative Agents(park_generative_2023, ) have been developed to enable agents to perform social actions. Additionally, AgentSim offers a comprehensive architecture for simulating standardized agent behaviors, incorporating a module for tool use among agents(lin_agentsims_2023, ).
@@
Nowadays, Researchers are gradually realizing the practical computational significance of an Agent’s personality on human-like behaviors(ghasem-aghaee_cognitive_2007, ; aher_using_2023, ; ma_understanding_2023, ; ziems_can_2024, ). In the study of Humanoid Agents(wang_humanoid_2023, ), researchers explored how emotions and needs influence an Agent’s planning, largely addressing the overly rational and stable machine-like state exhibited by former Agents(hagendorff_machine_2023, ). However, research on embodied agents in virtual environments seldom discusses other intrinsic dimensions. Expanding the scope to other Agent application scenarios, such as conversational interaction Chatbots(tu_characterchat_2023, ), we can still find some works focusing on the intrinsic aspects of Agents, such as employing psychological theories, especially cognitive psychology, to test and enhance the cognitive abilities of personality agents, and some works focusing on modelling emotions, especially emotional recognition and empathy(noauthor_emotional_nodate, ; habibi_empathetic_2023, ). In terms of character, some works simulate the static personality and knowledge of historical figures based on external data(noauthor_living_nodate, ), and others create chatbots with different static personalities based on personality psychology theories(tu_characterchat_2023, ).
@@
The intrinsic personality of real humans is not immutable. In psychology, researchers believe that personality is continuously influenced and changed by external factors(kampmann_feedback_2012, ). In existing discussions on research involving embodied agents, we rarely find studies detailing the process of such human-like intrinsic personality evolution or how intrinsic personality evolution affects an agent’s extrinsic behavior.